### v2

# Shaken, Not Stirred: A Novel Dataset for Visual Understanding of Glasses in Human-Robot Bartending Tasks

Abstract
Datasets for object detection often do not account for enough variety of glasses, due to their transparent and reflective properties. Specifically, open-vocabulary object detectors, widely used in embodied robotic agents, fail to distinguish subclasses of glasses. This scientific gap poses an issue to robotic applications that suffer from accumulating errors between detection, planning, and action execution. The paper introduces a novel method for the acquisition of real-world data from RGB-D sensors that minimizes human effort. We propose an auto-labeling pipeline that generates labels for all the acquired frames based on the depth measurements. We provide a novel real-world glass object dataset [^3] https://gajdosech2.github.io/GlassNICOLDataset/ that was collected on the Neuro-Inspired COLlaborator (NICOL), a humanoid robot platform. The data set consists of 7850 images recorded from five different cameras. We show that our trained baseline model outperforms state-of-the-art open-vocabulary approaches. In addition, we deploy our baseline model in an embodied agent approach to the NICOL platform, on which it achieves a success rate of 81% in a human-robot bartending scenario.
I INTRODUCTION
Transparent objects are everywhere, from households, healthcare assistance, and gastronomy to industrial and construction-site environments. Yet, these materials reflect only a very small fraction of visible light. Most of it is scattered and passes through. From a computer vision perspective, the processing of these objects is challenging but crucial for a successful deployment in real-world robotics applications [1]. As the intelligent robot starts to make its way into public spaces, where glass objects in the form of bottles and drinking glasses are an integral item, future generations of service robots are required to not only detect but also manipulate those materials. In addition, machine learning and data-driven methods require a large amount of training data to reach a suitable performance. Similar to robotic Sim2Real applications, existing research also proves the presence of a significant Sim2Real gap in the area of synthetic glass material generation [2, 3, 1], while those approaches require sophisticated photorealistic rendering methods.
Open‑vocabulary approaches are used for object detection in a plethora of embodied robotic setups, such as [4]. Embodied robotic agents, at a minimum, consist of a detection module, an action module, and an LLM utilized for high-level interaction and task planning. Recent approaches aim to fuse detection and language processing by utilizing vision-language models (VLMs) [5]. Being interactive by design, these agents are typically situated in scenarios that involve user interaction and often also collaboration, thus leaving a potentially infinite copora of object names to be detected.
Well-known open-vocabulary object detection approaches such as Grounding DINO [6], YOLOWorld [7], and OWL-Vit [8] have revolutionized object detection by expanding the range of recognizable object classes without extensive re‑annotation. These approaches are built upon end-to-end transformer‑based architectures. Global image features are extracted by decoder backbones and fused with language embeddings by utilizing deep metric learning. However, despite their success, our experiments suggest that these models exhibit a significant decrease in performance when encountering transparent objects, commonly found in HRI scenarios. To address these challenges, we introduce a novel real-world dataset that was captured on the humanoid robot NICOL shown in Fig. 1, including a new depth-based automated labeling method. The robot is equipped with a combination of RGB-D scanners and standard RGB cameras. Specific contributions of our work include:
  • A novel real-world dataset specifically designed for glass object detection, addressing the limitations of synthetic datasets and bridging the Sim2Real gap.
  • An automated labeling pipeline, shown in Fig. LABEL:fig:pipe, allowing a rapid annotation of data, leveraging existing visual foundation models.
  • A real-time integration of the visual detector with a physical robot, showing its potential in HRI scenarios.
  • A humanoid bartender task as a usecase, where the robot accurately detects glassware and executes pouring motions, establishing a platform for controlled experiments in realistic, socially engaging contexts.
II RELATED WORK
II-A Transparent Material Perception
Missing or incorrect RGB-D data for robotic perception can be estimated using implicit functions and geometric priors [9] and by balancing local and global depth features [10]. The depth map can also be de-projected into a point cloud and processed with techniques like 3D CNN to achieve depth completion [11]. Iterative refinement is also possible, with attention module on transparent regions [12] or with indirect geometry representations [13].
RGB often contains information missing from depth data, which can be leveraged with transfer learning [14]. Cues can be extracted from the RGB to fill-in depth observations using global optimization [1]. Incorporating affordance detection of the underlying object yield further improvement [15]. Other approaches include multimodality, such as visual-tactile fusion [16] and usage of polarization imaging [17].
Recently, neural radiance fields (NeRFs), have emerged as a promising tool for transparent object perception. GraspNeRF introduces a multiview-based 6-DoF grasp detection system [18]. Evo-NeRF extends this approach with real-time NeRF training and grasp adaptation [19]. Multi-view methods enhance transparent object perception by integrating multiple viewpoints [20]. Domain randomization techniques have also been employed to train models that generalize well to real-world scenarios [21]. The common limitation of existing approaches is their unpredictable performance on unseen scenes, highlighting the need for the methodology of automated benchmark data aggregation and annotation without human labor [2].
II-B Existing Datasets
The field of transparent object perception has long faced a scarcity of real-world datasets, with the majority of the work driven by synthetic data [2, 13, 22]. Synthetic datasets like Trans10k [23], SuperCaustics [24], and Dex-NeRF [25] offer a vast number of images (ranging from 9k to 100k) but suffer from the Sim2Real gap [26]. Besides common causes like the lack of noise and imperfections presented in real-world environments, transparent objects provide a unique challenge due to their complex light interaction. In contrast, real-world datasets like TransCG [27] and ClearGrasp [1] have attempted to address these limitations by collecting depth data using physical sensors. However, they lack diversity. Many datasets focus on isolated objects placed against controlled backgrounds, limiting their utility in highly cluttered environments [2]. The process of manual ground-truth depth and segmentation annotation remains a significant bottleneck [28]. Few works, like the Toronto Transparent Object Depth dataset leverage automated annotation techniques [11].
Furthermore, many existing large-scale RGB-D datasets [29] do not specifically focus on transparent objects [30]. We also note the existence of datasets like RGBP-Glass [31] and RGB-Thermal [32], introducing additional sensing modalities. Yet, these remain niche applications and do not fully address the generalization problem for robotic perception.
Figure 1: The NICOL humanoid robot used for down-stream task evaluation.
III Methodology
(a) Eye Camera RGB Texture
(b) Side RealSense RGB
(c) Fusion of all RGB-D Views
(d) Network Prediction
Figure 2: Different modalities present in our dataset, visualized on scene_145 from the validation split.
III-A NICOL Platform and Camera Model
The humanoid robot NICOL, shown in Fig. 1, has two 8-DoF arms that each have an anthropomorphic tendon-driven hand with 5-DoF attached. It has a 2-DoF 3D-printed head structure with an integrated face expression interface. The platform is designed for manipulation and HRI scenarios and offers a variety of sensors for different modalities. A tabletop of two meters in width and one meter in depth is attached in front of NICOL. The robot head contains two RGB fish-eye cameras with a 4k resolution. In our setup, an RGB-D Realsense D435if camera is mounted to NICOL’s head in addition, as well as two more RealSense cameras that are mounted to the left and right front pillars of the robot frame. Thus, the setup enables a complete depth perception of the objects on the table from multiple perspectives.
The moveable head and large area on the table covered by NICOL’s arms, combined with the fish-eye property of the RGB eyes, require a very precise camera calibration to allow for complex manipulation tasks such as pouring liquids. We use a semi-automated calibration method, which scans the table in front of NICOL from different angles to collect calibration data. Small calibration markers are attached to the tabletop that are automatically detected with a blob detection method. The detected positions in the camera image are manually mapped to the corresponding 3D ground truth positions by a human. The collected data is then used with the OpenCV camera calibration method [33] to calculate camera matrix $K$, the distortion coefficients $D$, as well as rotation matrix $R$ and the translation vector $t$. The calibration has to be performed individually for each of the three cameras used in our setup. We reached a reprojection error of $\sim 2~{}pixels$ for the RGB eye cameras and $\sim 4.5~{}pixels$ for the RGB-D camera. The two static RealSense cameras were calibrated with the well-known chessboard calibration method. Our camera model utilizes the camera parameters to provide a ray-casting method between the table and the camera sensor. In other words, pixels from one camera can be cast to the corresponding 3D coordinate at the tabletop and, for example back into one of the other cameras to mark a specific 3D coordinate in the images from all five cameras.
III-B Dataset Design
Recent approaches dealing with transparent object perception usually rely on synthetically rendered data [1, 3, 20]. The amount of real-world data from scenes containing these objects is severely limited, in comparison to datasets with objects from opaque materials [29, 2]. This is especially true for scenes with cluttered environments and a mix of transparent and non-transparent objects. We have designed a new benchmark dataset that is applicable to several computer vision and robotic tasks. Frames are captured from the egocentric view of the NICOL robot [34] performing various arm movements. Thus, complexity in various modalities is added as the dataset includes various degrees of occlusion, varying lighting conditions, and differing tabletop textures.
Specifically, our dataset consists of scenes with a mix of objects placed on a table in front of a humanoid robot. See Fig. 2 for an example scene. The collection process involves three separate passes:
  1. 1.
    Capturing the scene with clean glasses.
  2. 2.
    Placing 3D-printed green caps on top of the glasses for height measurement.
  3. 3.
    Replacing glasses with identical instances sprayed by chalk spray, making them opaque for obtaining depth ground truth geometry.
This design allows us to provide glass classification and detection labels, segmentation masks, as well as ground-truth depth measurements. Using the processing pipeline described in the following subsection, all annotations are created automatically, without a need for human labor.
Our setup consists of three RGB-D Realsense cameras, with one attached to the robot’s head and two at the sides of the table. Furthermore, there are two fisheye RGB cameras, one in each eye of the robot, as mentioned earlier. Using this setup, we captured 100 different scenes, with variations in color and texture of the table, lightning conditions, robot movements, and scene compositions. For each scene, there are 25 different views from the robot, acquired by a rotation in the neck joint. All cameras were calibrated and registered, allowing a projection of detection labels from the head-mounted RGB-D sensors to all other views.
In total, the sum of frames acquired from all five camera sensors (2 static RGB-D cameras and 25 images times 3 head cameras) results in a number of 77 frames per scene and 7700 frames in the training dataset. There are six different glass types present, each with its own class, namely: high beer glass, beer glass with a handle, wine glass, water glass, whisky glass, and shot glass. For each type, we also provide a ground-truth 3D model usable in surface reconstruction tasks [35]. Together with ground truth depth observations from the three RGB-D scanners, this can also be used in monocular depth estimation [36] and depth completion tasks [27]. Our main use case of the data presented in this paper is the detection and classification of a glass type. In a more general perception task, different glass types can be omitted, and all labels can be treated as general drinking glass annotations. Our data intentionally contained various unlabeled bottles. These provide negative examples in training for objects that are made from transparent material but would not be classified as any type of glass for drinking by a human.
In addition to the 7700 frames used for training and validation, we have also captured separate testing data. It contains 150 frames, with 75 frames captured using the same approach as the training data, plus another 75 frames captured during the bartender experiment described in Sec. IV-B, thus having a completely different character compared to the training data. Labels for the test data have been created manually, to mitigate the effect of label noise introduced by our automated pipeline described below.
III-C Auto-Labeling Pipeline
The labeling process uses depth sensing, color verification, and object detection to create accurate segmentation masks and bounding boxes. Initially, depth images are converted into a 3D point cloud, denoted as $P=\{p_{i}\in\mathbb{R}^{3}\}$ where each point $p_{i}=(x_{i},y_{i},z_{i})$ represents spatial coordinates obtained from the RealSense camera mounted on the robot’s head. The primary surface in the scene, corresponding to the table, is estimated using RANSAC-based plane fitting, yielding the plane parameters $a,b,c,d$. Points deviating significantly from this plane are identified as potential objects and subsequently clustered using a density-based clustering algorithm from Open3D [37], ensuring each cluster represents a discrete candidate object. Next, the height of the clusters is computed relative to the table plane using the formula:
$h=\frac{|d^{\prime}-d|}{\sqrt{a^{2}+b^{2}+c^{2}}}$ (1)
where $d^{\prime}$ is the cluster’s plane offset. We discard clusters with large height deviations from known glasses. The remaining candidates are projected onto the 2D image plane using intrinsic camera parameters, transforming 3D coordinates into pixel locations $(u,v)$ through perspective projection:
$\begin{bmatrix}u&v&1\end{bmatrix}^{T}\sim K\begin{bmatrix}x&y&z&1\end{bmatrix}^{T}$ (2)
where $K$ is the intrinsic calibration matrix of the camera. To further filter instances, color verification is performed in the CIELAB color space, ensuring that the detected object caps match the expected green hue. A final verification step utilizes YOLO-World, a deep-learning-based object detector, to confirm the detected object has glass-like features, mitigating false positives from other green objects in the scene [7]. As each of these steps work with different data modalities and different false-positive error profiles, we achieve a variant of a cascaded filtering approach by sensor fusion [38]. After obtaining the final candidates, points from the blobs are used as samples for the Segment Anything Model (SAM) [39], generating object masks for precise annotations. The segmentation masks $M_{i}$ are converted into polygons, defining the object’s shape in the image space. We derive bounding boxes $B_{i}=(x,y,w,h)$ from these masks, finalizing the object instance label. A visual overview of our labeling pipeline and an example annotation were presented earlier (see Fig. LABEL:fig:pipe).
III-D Baseline Detector
To demonstrate the advantages of our novel dataset, we use it to train an object detector and perform use-case scenarios and experiments. For real-time detection performance, we opted for an RTMDet architecture in its medium-size variant [40]. Naturally, different one-stage architectures, like the YOLO family [41], could be employed for this task.
The chosen network variant incorporates various training strategies. Specifically, the network employs a CSPNeXt backbone with a P5 architecture, a deepen factor of 0.67, and a widen factor of 0.75, enhanced with channel attention and synchronized batch normalization. The neck is a CSPNeXtPAFPN structure with two CSP blocks [42], facilitating multi-scale feature aggregation [40]. As proposed by the original authors, the head of the network calculates three loss functions. A Quality Focal Loss ($\beta=1.0$, loss weight = 1.0) for classification [43], Complete IoU (CIoU) Loss (loss weight = 2.0) for bounding box regression [44], and Dice Loss (loss weight = 3.0, reduction = “mean”) for mask prediction [40]. To increase robustness and grant better generalization for different scenarios, heavy data augmentation is employed. This includes techniques such as Cached Mosaic [45], Cached MixUp [46], Random Resize, Random Crop, and YOLOX-style HSV augmentation [47].
The training is performed from a pre-trained checkpoint [40], hence we only fine-tune it on our data. The learning rate is scheduled using a combination of Linear Warmup (start factor = 0.001, epochs 0–50) and MultiStepLR with decay at epochs 100, 200, and 400. The optimizer is AdamW [48] with a learning rate of 0.01 and a weight decay of 0.05. The model is trained for a total of 500 epochs with a batch size of 8, and evaluation is conducted every 50 epochs using the COCO metric for both bounding box and segmentation performance on 5 (representing a 5% split from our 100 scenes) validations scenes.
III-E Glass Base Points
To improve the robotic arm guidance without heavy modification of the detector architecture or lifting to 3D [49], we calculate 2D glass base positions separately. Given these locations and the known heights of the glasses from the classification, it enables robot pouring without exact 3D positions. We provide annotations for them in the dataset. By calculating the average position of each 3D-printed cap, we obtain point $\mathbf{p}=(x,y,z)^{\top}$ for each glass and collect them into a set $\mathbf{C}$. Taking the normal vector of the table $\mathbf{n}=(a,b,c)^{\top}$ and its normalized version $\mathbf{\hat{n}}$, the perpendicular distance to the table for each $\mathbf{p}\in\mathbf{C}$ is given by
$d_{\text{proj}}=\frac{ax+by+cz+d}{\|\mathbf{n}\|}.$ (3)
The projection of the point onto the table is computed as $\mathbf{p}_{\text{proj}}=\mathbf{p}-d_{\text{proj}}\mathbf{\hat{n}},$. To obtain 2D image coordinates, we use perspective projection with the camera matrix $K$. Taking 2D coordinates as centers of small bounding boxes with a fixed size, we introduce them as a separate keypoint class into the COCO object detection format.
Finally, we propose a modification of the detector head to obtain heat-map of glass base points. Before applying Non-Maximum Suppression [50], we extract all proposals of the keypoint class. We calculate the center location $(x_{c},y_{c})$ for each proposal and place Gaussian kernel $G(x,y)$ of size $k\times k$ (in our experiments $k=15$) at these locations:
$G(x,y)=\frac{1}{2\pi\sigma^{2}}\exp\left(-\frac{(x-x_{c})^{2}+(y-y_{c})^{2}}{2\sigma^{2}}\right).$ (4)
The contribution of each kernel is weighted by the confidence score $s_{i}$ of the corresponding bounding box. The final heatmap $H(x,y)$ is obtained by summing all contributions:
$H(x,y)=\sum_{i}s_{i}G_{i}(x,y)$ (5)
The 2D base points are extracted by identifying local maxima in the heatmap within each detected glass bounding box. See Fig. 2(d) for an example network output, with heatmap overlayed on glass detections. These points can also be calculated with the approach described in Section III-G. Full 3D localization using the depth maps provided in that dataset is proposed for future work.
Prompt 1: A summarized version of our system LLM prompt of the bartender task with an example of the pouring wine and pouring beer functionalities. You are deployed to function as a bartender. There are glass objects on the table. You should always respond as a bartender and offer available services like a bartender. When serving drinks, follow these rules:
1. Pouring Wine:
- Identify the correct bottle.
- Identify the correct glass.
- Example: $<$pour_wine(wine glass)$>$.
If there is only one wine glass, pour immediately into it without asking the user. If there are multiple wine glasses, ask the user which they prefer before pouring.
2. Pouring Beer:
- Identify the correct bottle.
- Identify the correct glass.
- Example: $<$pour_beer(beer glass)$>$.
- Example: $<$pour_beer(high beer glass)$>$.
If there is only one beer glass, pour immediately into it without asking the user. If there are multiple beer glasses, ask the user which they prefer before pouring.
You can fill a glass only once. When you pour into a glass, you should remember which glass is already filled. Always prioritize serving drinks accurately and safely. Respond to the user like a professional bartender.
III-F LLM Integration: Bartender Task
The NICOL platform supports an LLM integration for embodied agents in real-world robotic tasks [51]. We utilize the grounded LLM to implement a robot bartender scenario, leveraging the LLM’s reasoning to connect the robot’s sensory perception (identifying glass objects) and robot’s physical actions (precise arm manipulation for pouring). Our task facilitates natural interaction with users since the bartender scenario inherently involves social dynamics like recognizing drinkware, interpreting user requests, and responding appropriately. Although our implementation focuses on verbal communication, the task can be extended to incorporate context-driven and multimodal cues, including nonverbal communication. The LLM prompt is in Prompt III-E.
As shown in our previous work [4], the concept of procedural memory in LLMs is effective for rapid acquisition of text-based skills in robots, enabling them to perform adaptive and context-aware reasoning. We apply this concept to implement interactive use cases in the bartender task, where NICOL assists users by serving drinks. While not exhaustive, our scenario highlights the potential of integrating our proposed glass dataset and detector into a robotic application. The scenario consists of the following use cases:
  1. 1.
    Action-Object Alignment: The robot accurately maps the beverage pouring action to appropriate glassware in a way that aligns with user expectations and common conventions (e.g., pouring wine into a wine glass).
  2. 2.
    Resolving Ambiguities: The robot identifies potential ambiguities in object selection and seeks clarification through user corrective feedback. For example, if the user orders beer and multiple beer cups are available, the robot intuitively prompts the user to specify their preferred glass type, while providing a list of options.
  3. 3.
    Context-Driven Action Sequence: The robot uses LLM context memory to maintain continuity across the interaction, ensuring sequential tasks are carried out in logical order. For example, if the user orders a beer and then sequentially asks for another, then the robot can reason about which glass it already filled and choose the appropriate glass for the next pouring action.
III-G Robot Integration: Beverage Pouring Action
The Bézier curve-based motion planner described in our previous work [52] is utilized to design a pouring motion for the robot bartender task. The action picks up an already open bottle from a fixed position, moves it towards the opening of the glass, pours particles into it, and returns the bottle to its origin. Careful and precise pouring motions are required to prevent spilling particles and to avoid damaging the glasses [^1] No glass was harmed in any of our experiments.. The RealSense camera attached to the head of NICOL is used for this application, but it is also possible to use the left or right eye cameras. We use 3D-printed particles to imitate liquid as it prevents damage to the electrical hardware.
We use the projection approach of the camera model described in Sec. III-A to determine the glass coordinates in 3D space. First, the center of the bounding box bottom edge $b$ is projected from the image plane onto the table top. Thus, the approximately nearest point between the camera origin and the glass bottom is calculated. Since the given coordinate is located at the outer hull of the glass, an offset $o_{i}$ is added that takes the height and diameter of the i-th glass class into account. The center position of the glass opening changes its relation to the detected bounding box bottom center with respect to the glass’s position in the table plane and the height of the glass. The beverage pouring motion has to account for those spatial changes through a second dynamic pouring offset $p_{x}$ and $p_{y}$ that scales linearly with the specific x- and y-axis positions, as shown in Eq. 6:
$\begin{split}p_{x}=\epsilon\cdot p_{x}^{min}+(\epsilon\cdot\tau)p_{x}^{max}\\ p_{y}=\gamma\cdot p_{y}^{min}+(\gamma\cdot\tau)p_{y}^{max}\end{split}$ (6)
In the prior equation, suitable offsets for the smallest glass class $p_{x}^{min}$ and $p_{y}^{min}$ are adjusted with the linear scaling factor for the x-axis position $\epsilon$ and the y-axis position $\gamma$. In addition, suitable offsets for the highest glass class $p_{x}^{max}$ and $p_{y}^{max}$ are added by adjusting it with the linear scaling factor for the glass height $\tau$. As the height offset has shown to be influenced by the x- and y-position of the glass in our preliminary experiments, it is multiplied by the scaling factor for the corresponding axis $\epsilon$ or $\gamma$. The pouring offsets for the smallest and highest glass classes $p_{x}^{min}$, $p_{y}^{min}$, $p_{x}^{max}$, and $p_{y}^{max}$ can be determined by placing the smallest and highest glass at the maximum x- and y-position of the pouring workspace and manually tuning for the suitable offsets. The primitive scales well for the intermediate glass classes. The normalized scaling factors $\epsilon$, $\gamma$, and $\tau$ are calculated according to Eq. 7:
$\begin{split}\epsilon=&\frac{x_{n}-x_{min}}{x_{max}-x_{min}},\hskip 5.0pt\gamma=\frac{|y_{n}|-y_{min}}{y_{max}-y_{min}}\\ &\hskip 20.0pt\tau=\frac{h_{n}-h_{min}}{h_{max}-h_{min}}\\ \end{split}$ (7)
where the constants $x_{min}$, $x_{max}$, $y_{min}$, and $y_{max}$ are the workspace bounds. Similarly, $h_{min}$ and $h_{max}$ are the height of the smallest and highest glass class.
IV Experiments and Evaluation
We conduct two experiments to evaluate our system. First, we evaluate the performance of our fine-tuned RTMDet glass classifier using our proposed dataset and compare it against state-of-the-art off-the-shelf open-vocabulary object detectors. In the second experiment, we integrate our glass classifier with an LLM-powered embodied agent using the NICOL robot and perform an end-to-end system evaluation in a bartender-like Human-Robot scenario. In the next subsections, we show the evaluation metrics and results of each experiment.
IV-A Transparent Object Detection
We evaluate the detector on the test data using the two standard COCO evaluation metrics Average Precision and Average Recall at IoU from $0.5$ to $0.95$. Using data from the eye cameras in addition to the captures from the head-mounted RealSense during training slightly increases the performance. The values are calculated for two cases. In the first case (highlighted in Table I), we use the class names described in Sec. III-B as prompts to the open-vocabulary detectors. The shown values of AP and AR are the mean across all glass types (sometimes referred to as mAP and mAR, respectively). It should be mentioned that existing detectors are unaware of our classification of glass types. There is an ambiguity even for humans, such as the difference between a whisky and a water glass. Therefore, in the second case, we treat all instances as a general drink glass class. In preliminary experiments, we have also tried other prompts, such as transparent drink container, drinking glass, or drinkware with worse results. The testing split, as the rest of our data, does not contain any other objects classifiable as a drink glass by a human. We have also experimented with image-guided prompting with a template glass image. In such a configuration, the detectors preferred to look for other containers in a similar pose (such as from a top-right view), rather than focusing on the material and type. This caused an even higher false positive rate, with detections of objects like ceramic cups in a pose similar to the template.
The higher AP of our retrained model, even in the general drink glass setting, quantitatively proves our hypothesis that existing open-vocabulary detectors do not understand the concept of a drink container made from a glass material. In their latent space, it seems to be confused with containers made from arbitrary different materials, such as a metal cans. They also cannot distinguish between transparent containers like bottles and drinking glasses, whereas for humans they serve different purposes. Going one additional step towards glass type classification makes existing detectors basically unusable for our scenario. In terms of AR, existing models perform more appropriately. They do not have many false negatives, generally selecting everything that looks like a rounded container.
TABLE I: Comparison of glass detection using zero-shot detectors and transfer-learned lightweight model
Method AP@[0.5:0.95] AR@[0.5:0.95]
general
class
glass
types
general
class
glass
types
OWL-Vit[8] 0.403 0.014 0.671 0.112
G-DINO[6] 0.638 0.102 0.829 0.357
YOLO-World[7] 0.707 0.191 0.844 0.413
*RTMDet-M[40]
(head RGB only)
0.774 0.713 0.850 0.801
*RTMDet-M[40]
(all RGB views)
0.786 0.718 0.854 0.809
* Our fine-tuned models.
IV-B Robot Integration: Bartender Task
In this experiment, we use the LLM integration (see Sec. III-F) and beverage pouring motion (see Sec. III-G) to evaluate the performance of our robot bartender task, deployed on the physical robot hardware. We evaluate our system in an end-to-end manner, i.e., a user (an experimenter), interacts with the system in a way that resembles a real-world scenario. Our experimental setup consists of NICOL with six glass objects placed on the table – a shot glass, wine glass, whiskey glass, water glass, beer glass, and high beer glass. Additionally, two plastic bottles – a wine bottle and beer bottle – are placed on each side of the table and filled with artificial liquid made from 3D-printed round-shaped particles with distinct colors. The pouring workspace has a width of 55 cm and a depth of 35 cm. The positions of the bottles were swapped halfway through the experiment for a fair comparison of the pouring by each arm. The user stands across the table and issues verbal commands through an external microphone, processed via a Whisper integration.
Each interaction consists of the following: the user orders wine, then a beer, followed by another beer request. We select these interactions to assess the system’s ability to adapt to different beverage types, resolve ambiguities in glassware selection, and ensure a logical action sequence. This can be noted as a three-phase process: 1)  User Order Recognition: The system processes the user’s order and determines the appropriate drink and glassware, through action-object mapping or user preference. For example: if the user says “I’d like a wine”, the robot must infer that wine should be poured into a wine glass, 2) Reasoning & Decision Making: Based on user input, the robot reasons about the correct bottle and glass object. Any ambiguity is clarified through user feedback. For example: if the user says “Can I get a beer for my friend?”, the robot must decide the correct glassware based on context and prior interactions (which glass is already filled), 3) Action Execution: The robot performs a pouring action correctly to serve the requested drink. We use OpenAI GPT‑4o mini and reset the chat after each interaction.
We conducted 35 trials, meaning that 105 pouring actions were executed in total. We use the Success rate of the end-to-end pouring action as a metric to evaluate our experiment. However, we report the average error rate in the following categories: Spill referring to slight spilling of particles outside the intended glass, LLM representing LLM reasoning inaccuracies, and Detection, i.e., errors in the glass detector (cf. Fig. 3). Each pouring motion is considered successful if the robot correctly pours the target drink into the target glass without any particles spilled. Our system achieves a success rate of 81% in the pouring action across the two robot arms, while only 4.8% of the cases exhibited minor spills of particles outside the designated glass, demonstrating the suitability of our approach for precise pouring in a bartender task. Our LLM integration showed high resilience in generating correct robot actions across the different use cases: action-object mapping, ambiguity resolution, and logical action sequencing. In 10.5% of the cases, user correction was needed due to reasoning inconsistencies, such as suggesting the wrong glass for a specific drink. However, only one instance was recorded, where the LLM generated an unintended action, mistakenly handing over the object instead of pouring into it, leading to a failure case. Only 4.8% of the cases resulted in failures due to occasional glass detector errors, such as failing to detect the beer glass.
Figure 3: Results of our end-to-end evaluation of the NICOL robot bartender task, highlighting success rates, spilling occurrences, errors in LLM reasoning, and glass detection.
V Discussion
As can be seen in Table I, our new benchmark dataset helps to push the current state-of-the-art detector boundaries. Zero-shot models like YOLO-World[7], OWL-Vit [8], and GroundingDINO[6] are widely used for opaque models, but their performance deteriorates on concepts of transparency, glass material and drink-purposed containers. On the other hand, our processing pipeline allows us to train a lightweight model from RTMDet [40] family, achieving a high precision with a smaller rate of false positives. This proves the fact that reliable recognition of glass objects is far from achieved with general models. The glass detection system has shown high performance when integrated with the physical robot, achieving an overall success rate of 81% in a bartending scenario. While the results show that there is still room for improvements in our object detection and robotic action modules, the biggest error source originates from the LLM utilized for high-level planning, making more than 50% of the failure trials. Unfortunately, the glass base keypoints described in Sec. III-E oscillated too much to be applicable in the bartending agent in preliminary experiments. However, 3D planning approaches seem to be a promising area for future research. Our system already shows high performance by planning the motion based on the 2D image perception of the scene. A 3D perception approach, e.g., by utilizing the chalk samples delivered with our dataset for depth estimation, would thus still be capable of improving it.
VI CONCLUSION
This paper fills a critical gap in glass object perception by providing a novel real-world dataset. The proposed auto-labeling method reduces the human effort for data aggregation in similar tasks to an absolute minimum. We contribute three distinct repositories under an open-source license: the glass dataset, the glass detector model and experiment code, and the NICOL camera model. The quality of our dataset and camera model allows for very precise projections between different camera perspectives. Our glass detector outperforms SOTA open-vocabulary approaches and reaches a baseline success rate of 81% when integrated with our embodied bartending agent. Future work will include expanding the dataset with more diverse scenes and improving robotic manipulation strategies through enhanced LLM integration. The most interesting next step for the action module is a proprioceptive 3D planning approach that can dynamically adapt to spatial changes, like a user moving a glass while pouring, fostering seamless Human-Robot collaboration in real-world tasks.
Acknowledgements
We thank Svorad Štolc from Photoneo for external consultations of the vision part, Philipp Allgeuer for his contribution to the NICOL robot, and Connor Gäde for his contribution to the bartending experiment.